# Noise-Domain Non-Orthogonal Multiple Access

### v1

@@ Abstract
In this paper, we present noise-domain non-orthogonal multiple access (ND-NOMA), an innovative communication scheme that utilizes the modulation of artificial noise mean and variance to convey information. Distinct from traditional methods such as power-domain non-orthogonal multiple access (PD-NOMA) that heavily relies on successive interference cancellation (SIC), ND-NOMA utilizes the noise domain, considerably reducing power consumption and system complexity. Inspired by noise modulation, ND-NOMA enhances energy efficiency and provides lower bit error probability (BEP), making it highly suitable for next-generation Internet-of-things (IoT) networks. Our theoretical analyses and computer simulations reveal that ND-NOMA can achieve exceptionally low bit error rates in both uplink and downlink scenarios, in the presence of Rician fading channels. The proposed multi-user system is supported by a minimum distance detector for mean detection and a threshold-based detector for variance detection, ensuring robust communication in low-power environments. By leveraging the inherent properties of noise, ND-NOMA offers a promising platform for long-term deployments of low-cost and low-complexity devices.
Index Terms:
@@
I Introduction
Orthogonal multiple access (OMA) techniques offer advantages such as effective interference management, simplicity of implementation, and fair resource allocation. However, they also present several disadvantages, including reduced efficiency as number of users increases, challenges in dynamic resource allocation, scalability issues in large-scale networks, and limited flexibility due to the need to maintain orthogonality among users [1, 2, 3]. While OMA is well-suited for current wireless networks, future scenarios with high-density and dynamic demands may benefit from newer techniques like non-orthogonal multiple access (NOMA).
@@ I Introduction
However, NOMA systems face several challenges, such as optimizing power allocation to balance energy efficiency and quality-of-service, requiring accurate channel state information (CSI) for effective SIC operation, and managing residual interference from imperfect SIC. Additionally, there are trade-offs between spectral and energy efficiency, receiver design complexities, SIC error propagation, and sensitivity to channel gain measurements [6]. Interference management, integration with carrier aggregation, and addressing security concerns are also critical for NOMA systems [7].
In earlier times, thermal noise modulation (TherMod) was introduced to convey information using thermal noise [8]. This concept recently evolved into schemes that transmit data using parameters like random signal variance. A transceiver architecture later emerged, using either artificially generated or thermal noise sources [9]. Similarly, noise loop modulation had been utilized to deliver artificial noise in a feedback loop between legitimate users, enhancing unconditional security by preventing eavesdropping without requiring knowledge of the eavesdropper’s channel [10]. Parallel to these studies, during transmission variance based sampling was also previously used[11]. It has been shown that noise modulation might provide advantages for simplifying the receiver architectures and robustness to certain channel effects and system impairments. This motivates to consider its application in multi-user environments.
@@
In this work, by taking the noise modulation concept one step further, we propose an innovative communication scheme called noise-domain NOMA (ND-NOMA). This scheme utilizes the mean and variance of the artificial noise signals to transmit data. Basically, bits of one user are modulated using the mean of the noise samples, while bits of the other user are modulated using the variance of the noise samples. Since ND-NOMA leverages noise samples, the system does not require processing steps such as time synchronization and equalization for the downlink scenario. Thus, ND-NOMA significantly reduces power consumption and system complexity, making it ideal for the low power consumption requirements of next-generation applications such as Internet-of-things (IoT) networks. Our theoretical analyse and computer simulations demonstrate that ND-NOMA can achieve exceptionally low bit error rate (BER) in both uplink and downlink scenarios. The robust performance of the system is supported by a minimum distance detector for mean detection and a threshold-based detector for variance detection. and ensuring reliable communication even in low-power scenarios. Our findings indicate that ND-NOMA has excellent performance particularly in Rician fading channels and can be utilized for long-term and large-scale deployments of low-cost, low-complex devices.
ND-NOMA is well-suited for integration with Device B types of ambient IoT systems, as highlighted in [12]. Device B, with its limited energy storage and semi-passive operation, aligns with ND-NOMA’s low-power and low-complexity demands, essential for sustainable IoT deployments. By utilizing noise mean and variance for data transmission, ND-NOMA reduces the need for power-intensive processes like synchronization, making it ideal for energy-efficient systems. Its robustness in low-power scenarios further enhances the performance of ambient IoT networks, positioning ND-NOMA as a strong contender for next-generation applications. Studies in ISAC show that random waveforms enhance spectral efficiency, complementing ND-NOMA’s goals of low-power communication [13]. Moreover, energy harvesting techniques such as SWIPT further support ND-NOMA’s energy-efficient design [14, 15].
The fascinating property of Gaussian distribution is that the sum of two independent Gaussian variables remains Gaussian [16]. In ND-NOMA, we exploit this by adding two Gaussian noise signals, each carrying different user information. The resulting signal retains its Gaussian nature, simplifying processing and detection at the receiver. This allows ND-NOMA to efficiently handle multiple signals with minimal error, ideal for low-power IoT networks.
@@
II Uplink ND-NOMA: System Model and Performance Analysis
@@
II-A System Model
As shown in Fig. 1, in the uplink transmission of two users, one of the users (User 1, $\text{U}_{1}$) uses the mean of the transmitted Gaussian samples. In contrast, the other user (User 2, $\text{U}_{2}$) exploits the variance as in noise modulation schemes. Specifically, denoting the $n$th noise samples of $\text{U}_{1}$ and $\text{U}_{2}$ respectively by $s_{1}^{n}$ and $s_{2}^{n}$, for $\text{U}_{1}$, the transmission of information bits is accomplished by alternating the mean of the samples between low and high values, that is, for bit-00 and bit-$1$, we have, $s_{1}^{n}\sim\mathcal{N}(m_{1,l},\sigma_{1}^{2})$ and $s_{1}^{n}\sim\mathcal{N}(m_{1,h},\sigma_{1}^{2})$. As discussed later, we set $m_{1,h}=-m_{1,l}$ for optimum performance. We assume that two users consider $N$ noise samples for each bit.
On the other hand, for $\text{U}_{2}$, we apply noise modulation by alternating the noise variance only, and for bit-00 and bit-$1$, we have $s_{2}^{n}\sim\mathcal{N}(m_{2},\sigma_{2,l}^{2})$ and $s_{2}^{n}\sim\mathcal{N}(m_{2},\sigma_{2,h}^{2})$. At this point, we set the mean of samples to zero, that is, $m_{2}=0$, to minimize the interference to $\text{U}_{1}$. This initial model assumes binary-level signaling, while a generalization is subject to further studies.
In this setup, we assume an average transmission power of $P$ for each user, which corresponds to the second moment of the transmitted samples [^1] Assuming equiprobable $\text{U}_{2}$ bits, over a long sequence, half of its samples will have $\sigma_{2,l}^{2}$ variance while the other half having $\sigma_{2,h}^{2}$. As a request, the second moment of $\text{U}_{2}$ samples converges to $P$.: $E[(s_{1}^{n})^{2}]=E[(s_{2}^{n})^{2}]=P$, which equals to the sum of squared mean and variance of their samples. In light of this, $P=m_{1,l}^{2}+\sigma_{1}^{2}$ for $\text{U}_{1}$. Here, $\text{U}_{1}$ dedicates $\beta$ portion of its available power for the variance of its samples while dedicating a $(1-\beta)$ portion for the mean, that is, $\sigma_{1}^{2}=\beta P$ and $m_{1,l}^{2}=m_{1,h}^{2}=(1-\beta)P$. Since $m_{2}=0$, all power of $\text{U}_{2}$ is dedicated to the variance of its samples, $P=(\sigma_{2,l}^{2}+\sigma_{2,h}^{2})/2$.
@@ II-A System Model
Here, $h_{1}$ and $h_{2}$ are the complex baseband channel coefficients of the links between $\text{U}_{1}$-BS and $\text{U}_{2}$-BS, respectively, and $w^{n}\sim\mathcal{CN}(0,\sigma_{w}^{2})$ is the complex baseband additive white Gaussian noise sample at the BS. Our analysis and computer simulations assume either unit-gain Rayleigh or Rician fading models. Accordingly, for Rayleigh fading, we have $h_{1},h_{2}\sim\mathcal{CN}(0,1)$ while for Rician fading with $K$ parameter, we have $h_{1,R},h_{1,I},h_{2,R},h_{2,I}\sim\mathcal{N}\left(\sqrt{\frac{K}{2(1+K)}},\frac{1}{2(1+K)}\right)$. Then, conditioned on channel coefficients, we obtain
@@
II-B Uplink - User 1 Detection
@@
The above detector is a “minimum distance detector” and considering the complex Gaussian distribution of $\bar{y}$, it is a maximum likelihood one. In light of this and also considering the symmetry for bit-00 and bit-$1$, conditioned on channel coefficients, the BEP of $\text{U}_{1}$ becomes
@@
Finally, we note that, as NoiseMod schemes, conditional BEP is a decaying function of $\sqrt{N}$ (through $\sigma_{D}^{2}$). Section IV investigates the unconditional BEP over different types of fading channels.
II-C Uplink - User 2 Detection
In contrast to $\text{U}_{1}$’s mean detection problem, the task of the BS is to formulate a variance detection problem to extract $\text{U}_{2}$’s bits. While we work on a real and positive search space for this problem, complex distributions of $y^{n}$ and $\bar{y}$, more importantly, their correlated real and imaginary components, possess a severe challenge on the BEP derivation of $\text{U}_{2}$.
@@ II-C Uplink - User 2 Detection
Here, $\eta$, $\delta$, and $\alpha$ are system-related constants defined in Section II.A. Defining a variance threshold as $\gamma=\chi\sigma_{w}^{2}$, $\text{U}_{2}$ detection problem is formulated as
@@
where $\mu_{s}$ and $\sigma_{s}$ are conditioned on the transmitted bit through $\sigma_{2,k}^{2}$. Selecting the threshold for equal bit error probabilities of bit 00 and bit $1$ yields
@@
Despite its simple detection architecture and our assumptions above, the final result for the BEP is still too complicated for taking an average over channel realizations, and we have to use numerical integration as in (9) to obtain unconditional BEP.
III Downlink ND-NOMA: System Model and Performance Analysis
In this section, we emphasize downlink ND-NOMA for two users by defining its system model and then presenting its theoretical BEP performance.
TABLE I: Statistics of the transmitted and received signals for Downlink ND-NOMA
@@ III Downlink ND-NOMA: System Model and Performance Analysis
III-A System Model
As shown in 2, for the downlink scheme, a composite signal is generated by the BS considering the information bits of two users. Here, we again consider binary modulation for simplicity. Accordingly, BS’s transmitted noise samples follow $s_{BS}^{n}\sim\mathcal{N}(m_{1,i},\sigma_{2,k}^{2})$ distribution for $i,k\in\left\{l,h\right\}$. In other words, the mean of the transmitted signal is dictated by $\text{U}_{1}$’s bit while its variance is determined according to $\text{U}_{2}$’s bit. In this case, the received signals at two users are given by
@@ III-A System Model
for $n=1,\ldots,N$. Here, $h_{1}$ and $h_{2}$ stand for the downlink channel fading coefficients between the BS and users, and $w_{1}^{n}$ and $w_{2}^{n}$ are AWGN samples at users with $\mathcal{CN}(0,\sigma_{w}^{2})$ distribution. Specifically, Table I lists the distribution of the transmitted and received signals at all terminals depending on four user bit combinations, $00,01,10$, and $11$, where the first and second bits respectively stand for $\text{U}_{1}$ and $\text{U}_{2}$ bits for a given bit duration.
Here, we assume that the total transmission power is fixed to $P$, that is, $E[(s_{BS}^{n})^{2}]=P$. Similar to the $\beta$ parameter of the uplink scheme, we define a new parameter, $\psi$, as the portion of $\text{U}_{1}$’s allocated power, that is, $m_{1,l}^{2}=m_{1,h}^{2}=\psi P$ (dc power of the transmitted signal). Accordingly, for $\text{U}_{2}$, we have $(\sigma_{2,l}^{2}+\sigma_{2,h}^{2})/2=(1-\psi)P$ (ac power of the transmitted signal). We again consider $m_{1,h}=-m_{1,l}$ and $\sigma_{2,h}^{2}=\alpha\sigma_{2,l}^{2}$. Later, we will show that the performance of $\text{U}_{1}$ and $\text{U}_{2}$ highly depends on the selection of $\psi$ parameter.
@@
III-B Downlink - User 1 Detection
The receiver architecture of the downlink scheme for $\text{U}_{1}$ is very similar to that of the uplink scheme, and the BEP can be evaluated by simple modifications. Specifically, $\text{U}_{1}$ calculates the sample mean of its received samples as $\bar{y}_{1}=\frac{1}{N}\sum\nolimits_{n=1}^{N}y^{n}_{1}$, which is also Gaussian distributed with $\mathrm{E}[\bar{y}_{1}]=h_{1}m_{1,i}$ and $\mathrm{VAR}[\bar{y}_{1}]=(|h_{1}|^{2}\sigma_{2,k}^{2}+G_{w}^{2})/N$ for $i,k\in\{l,h\}$. Accordingly, the minimum distance detector is formulated as

### v2

# IEEEexample:BSTcontrol

@@ Abstract
In this paper, we present noise-domain non-orthogonal multiple access (ND-NOMA), an innovative communication scheme that utilizes the modulation of artificial noise mean and variance to convey information. Distinct from traditional methods such as power-domain non-orthogonal multiple access (PD-NOMA) that heavily rely on successive interference cancellation (SIC), ND-NOMA utilizes the noise domain, considerably reducing power consumption and system complexity. Inspired by noise modulation, ND-NOMA provides lower bit error probability (BEP), making it highly suitable for next-generation Internet-of-things (IoT) networks. Our theoretical analyses and computer simulations reveal that ND-NOMA can achieve exceptionally low bit error rates in both uplink and downlink scenarios, in the presence of Rician fading channels. The proposed multi-user system is supported by a minimum distance detector for mean detection and a threshold-based detector for variance detection, ensuring robust communication in low-power environments. By leveraging the inherent properties of noise, ND-NOMA offers a promising platform for long-term deployments of low-cost and low-complexity devices.
{IEEEkeywords}
@@
1 INTRODUCTION
[…]
Orthogonal multiple access (OMA) techniques offer advantages such as effective interference management, simplicity of implementation, and fair resource allocation. However, they also present several disadvantages, including reduced efficiency as the number of users increases, challenges in dynamic resource allocation, scalability issues in large-scale networks, and limited flexibility due to the need to maintain orthogonality among users [1, 2, 3]. While OMA is well-suited for current wireless networks, future scenarios with high-density and dynamic demands may benefit from newer techniques like non-orthogonal multiple access (NOMA).
@@ 1 INTRODUCTION
Figure 1: Illustration of a potential ND-NOMA use-case in low-data-rate, low-energy IoT environments, where (a) represents the downlink transmission and (b) represents the uplink transmission. For illustration, the legacy user is depicted as a sinusoidal carrier, whereas ND-NOMA users exhibit a noise-based waveform.
However, NOMA systems face several challenges, such as optimizing power allocation to balance energy efficiency and quality-of-service, requiring accurate channel state information (CSI) for effective SIC operation, and managing residual interference from imperfect SIC. Additionally, there are trade-offs between spectral and energy efficiency, receiver design complexities, SIC error propagation, and sensitivity to channel gain measurements [6]. Recent studies have also explored enhancing the physical layer security of uplink NOMA systems through the use of energy-harvesting jammers, offering a promising direction for secure and energy-aware NOMA designs [7]. Interference management, integration with carrier aggregation, and addressing security concerns are also critical for NOMA systems [8].
In earlier times, thermal noise modulation (TherMod) was introduced to convey information using thermal noise [9]. This concept recently evolved into schemes that transmit data using parameters like random signal variance. A transceiver architecture later emerged, using either artificially generated or thermal noise sources [10]. Similarly, noise loop modulation had been utilized to deliver artificial noise in a feedback loop between legitimate users, enhancing unconditional security by preventing eavesdropping without requiring knowledge of the eavesdropper’s channel [11]. Parallel to these studies, during transmission variance-based sampling was also previously used[12]. It has been shown that noise modulation might provide advantages for simplifying the receiver architectures and robustness to certain channel effects and system impairments. This motivates the consideration of its application in multi-user environments.
@@
In this work, by taking the noise modulation concept one step further, we propose an innovative communication scheme called noise-domain NOMA (ND-NOMA). The fascinating property of the Gaussian distribution is that the sum of two independent Gaussian variables remains Gaussian [13]. In ND-NOMA, we exploit this by adding two Gaussian noise signals, each carrying different user information. The resulting signal retains its Gaussian nature, simplifying processing and detection at the receiver, which allows the ND-NOMA to efficiently handle multiple signals with minimal error, ideal for low-power IoT networks. This scheme utilizes the mean and variance of the artificial noise signals to transmit data. Basically, bits of one user are modulated using the mean of the noise samples, while bits of the other user are modulated using the variance of the noise samples. Since ND-NOMA leverages Gaussian noise samples, the system does not require processing steps such as frequency synchronization for the downlink scenario. Thus, ND-NOMA significantly reduces power consumption and system complexity, making it ideal for the low power consumption requirements of next-generation applications such as Internet-of-things (IoT) networks. Our theoretical analysis and computer simulations demonstrate that ND-NOMA can achieve exceptionally low bit error rate (BER) in both uplink and downlink scenarios. The robust performance of the system is supported by a minimum distance detector for mean detection and a threshold-based detector for variance detection, ensuring reliable communication even in low-power scenarios. Our findings indicate that ND-NOMA has excellent performance, particularly in Rician fading channels, and can be utilized for long-term and large-scale deployments of low-cost, low-complex devices.
In emerging IoT use-cases, legacy systems were not performing well enough in terms of scalability, so the low-data-rate communication schemes were essential. Unlike PD-NOMA or CD-NOMA, designed for high-throughput scenarios, ND-NOMA offers a noise-domain access strategy well suited for low-data-rate IoT environments. As next-generation multiple access technology is a key candidate for 6G, ND-NOMA aligns well with the access needs of future wireless communication scenarios. ND-NOMA is well-suited for integration with Device B types of ambient IoT systems, as highlighted in [14]. Device B, with its limited energy storage and semi-passive operation, aligns with ND-NOMA’s low-power and low-complexity demands, essential for sustainable IoT deployments. For Device Type B, the goal is to have low complexity, and with its simple receiver architecture, ND-NOMA is well-suited for low-data-rate scenarios (0.1–5 kbps). Device B is semi-passive, uses backscattering with limited energy storage, and has power consumption below 1 mW. ND-NOMA removes the need for SIC and strict frequency and phase synchronization, reducing computational and energy costs at receivers, and making it ideal for large-scale IoT, thereby further lowering complexity and power consumption for practical deployments. Studies in ISAC show that random waveforms enhance spectral efficiency, complementing ND-NOMA’s goals of low-power communication [15]. Moreover, energy harvesting techniques such as SWIPT further support ND-NOMA’s energy-efficient design [16, 17]. It is worth mentioning that the ND-NOMA scheme is not designed for broadband cellular communication scenarios, but instead it is designed for low-power and low-data-rate IoT environments, and our main motivation is to decompose the signals without using SIC. Using noise statistics such as mean and variance to convey information requires a sufficient number of noise samples per symbol for accurate estimation. Without increasing the symbol duration or the sampling frequency, estimation accuracy may degrade, particularly in fast-varying channels. Fig. 1 provides a generic representation of the ND-NOMA concept within the context of IoT applications; however, the specific operational mechanisms and signal structures are more clearly illustrated in Figs. 2 and 3. ND-NOMA suits low-data-rate smart city applications such as environmental monitoring, smart metering, and public infrastructure telemetry, where energy efficiency, scalability, and asynchronous transmission are essential at the receivers.
@@
2 Uplink ND-NOMA: System Model and Performance Analysis
@@
2.1 System Model
As shown in Fig. 2, in the uplink transmission of two users, one of the users (User 1, $\text{U}_{1}$) uses the mean of the transmitted Gaussian samples. In contrast, the other user (User 2, $\text{U}_{2}$) exploits the variance as in noise modulation schemes. Specifically, denoting the $n$th noise samples of $\text{U}_{1}$ and $\text{U}_{2}$ respectively by $s_{1}^{n}$ and $s_{2}^{n}$, for $\text{U}_{1}$, the transmission of information bits is accomplished by alternating the mean of the samples between low and high values, that is, for bit-0 and bit-$1$, we have, $s_{1}^{n}\sim\mathcal{N}(m_{1,l},\sigma_{1}^{2})$ and $s_{2}^{n}\sim\mathcal{N}(m_{1,h},\sigma_{1}^{2})$. As discussed later, we set $m_{1,h}=-m_{1,l}$ for optimum performance. We assume that two users consider $N$ noise samples for each bit.
On the other hand, for $\text{U}_{2}$, we apply noise modulation by alternating the noise variance only, and for bit-0 and bit-$1$, we have $s_{2}^{n}\sim\mathcal{N}(m_{2},\sigma_{2,l}^{2})$ and $s_{2}^{n}\sim\mathcal{N}(m_{2},\sigma_{2,h}^{2})$. At this point, we set the mean of samples to zero, that is, $m_{2}=0$, to minimize the interference to $\text{U}_{1}$. This initial model assumes binary-level signaling, while a generalization is subject to further studies.
In this setup, we assume an average transmission power of $P$ for each user, which corresponds to the second moment of the transmitted samples [^1] Assuming equiprobable $\text{U}_{2}$ bits, over a long sequence, half of its samples will have $\sigma_{2,l}^{2}$ variance while the other half having $\sigma_{2,h}^{2}$. As a result, the second moment of $\text{U}_{2}$ samples converges to $P$.: $E[(s_{1}^{n})^{2}]=E[(s_{2}^{n})^{2}]=P$, which equals the sum of the squared mean and variance of their samples. In light of this, $P=m_{1,l}^{2}+\sigma_{1}^{2}$ for $\text{U}_{1}$. Here, $\text{U}_{1}$ dedicates $\beta$ portion of its available power for the variance of its samples while dedicating a $(1-\beta)$ portion for the mean, that is, $\sigma_{1}^{2}=\beta P$ and $m_{1,l}^{2}=m_{1,h}^{2}=(1-\beta)P$. Since $m_{2}=0$, all power of $\text{U}_{2}$ is dedicated to the variance of its samples, $P=(\sigma_{2,l}^{2}+\sigma_{2,h}^{2})/2$.
@@ 2.1 System Model
Here, $h_{1}$ and $h_{2}$ are the complex baseband channel coefficients of the links between $\text{U}_{1}$-base station (BS) and $\text{U}_{2}$-BS, respectively, and $w^{n}\sim\mathcal{CN}(0,\sigma_{w}^{2})$ is the complex baseband additive white Gaussian noise sample at the BS. Our analysis and computer simulations assume either unit-gain Rayleigh or Rician fading models. Accordingly, for Rayleigh fading, we have $h_{1},h_{2}\sim\mathcal{CN}(0,1)$ while for Rician fading with $K$ parameter, we have $h_{1,R},h_{1,I},h_{2,R},h_{2,I}\sim\mathcal{N}\left(\sqrt{\frac{K}{2(1+K)}},\frac{1}{2(1+K)}\right)$. Then, conditioned on channel coefficients, we obtain
@@
2.2 Uplink - User 1 Detection
@@
The above detector is a “minimum distance detector” and considering the complex Gaussian distribution of $\bar{y}$, it is a maximum likelihood one. In light of this and also considering the symmetry for bit-0 and bit-$1$, conditioned on channel coefficients, the BEP of $\text{U}_{1}$ becomes
@@
Finally, we note that, as NoiseMod schemes, conditional BEP is a decaying function of $\sqrt{N}$ (through $\sigma_{D}^{2}$).
2.3 Uplink - User 2 Detection
In contrast to $\text{U}_{1}$’s mean detection problem, the task of the BS is to formulate a variance detection problem to extract $\text{U}_{2}$’s bits. While we work on a real and positive search space for this problem, complex distributions of $y^{n}$ and $\bar{y}$, more importantly, their correlated real and imaginary components, pose a severe challenge to the BEP derivation of $\text{U}_{2}$.
@@ 2.3 Uplink - User 2 Detection
Here, $\eta$, $\delta$, and $\alpha$ are system-related constants defined in Section II.A. Defining a variance threshold as $\gamma=\chi\sigma_{w}^{2}$, where $\chi$ is the scaled threshold value [10], and it is defined as
$\chi=\frac{2(1+\delta)(1+\alpha\delta)}{2+\delta(1+\alpha)}.$ (12)
Thus, $\text{U}_{2}$ detection problem is formulated as
@@
where $\mu_{s}$ and $\sigma_{s}$ are conditioned on the transmitted bit through $\sigma_{2,k}^{2}$. Selecting the threshold for equal bit error probabilities of bit 0 and bit $1$ yields
@@
Despite its simple detection architecture and our assumptions above, the final result for the BEP is still too complicated for taking an average over channel realizations, and we have to use numerical integration as in (9) to obtain unconditional BEP. Nevertheless, the computational complexity of our approach remains significantly lower than traditional PAM demodulation methods with SIC, which require iterative interference cancellation and complex computations. In contrast, our detection methods rely solely on straightforward mean, variance, and threshold calculations, making them far more efficient and less computationally demanding.
The following steps outline the key aspects of the uplink system model, summarizing the transmission and detection processes:
  • Step 1: U1 encodes bits by adjusting the mean of its Gaussian noise signal, while U2 encodes bits by varying the variance with its mean kept at zero to avoid interference.
  • Step 2: The base station receives the combined signals from U1 and U2, affected by channel fading and noise.
  • Step 3: The base station decodes U1’s bit by computing the mean of received samples and uses a minimum distance detector, while U2’s bit is decoded by computing the variance and applying a threshold-based detector.
3 Downlink ND-NOMA: System Model and Performance Analysis
In this section, we focus on downlink ND-NOMA for two users by defining its system model and then presenting its theoretical BEP performance.
Table 1: Statistics of the transmitted and received signals for Downlink ND-NOMA
@@ 3 Downlink ND-NOMA: System Model and Performance Analysis
3.1 System Model
As shown in Fig. 3, for the downlink scheme, a composite signal is generated by the BS considering the information bits of two users. Here, we again consider binary modulation for simplicity. Accordingly, BS’s transmitted noise samples follow $s_{BS}^{n}\sim\mathcal{N}(m_{1,i},\sigma_{2,k}^{2})$ distribution for $i,k\in\left\{l,h\right\}$. In other words, the mean of the transmitted signal is dictated by $\text{U}_{1}$’s bit while its variance is determined according to $\text{U}_{2}$’s bit. In this case, the received signals at two users are given by
@@ 3.1 System Model
for $n=1,\ldots,N$. Here, $h_{1}$ and $h_{2}$ stand for the downlink channel fading coefficients between the BS and users, and $w_{1}^{n}$ and $w_{2}^{n}$ are AWGN samples at users with $\mathcal{CN}(0,\sigma_{w}^{2})$ distribution. Specifically, Table 1 lists the distribution of the transmitted and received signals at all terminals depending on four user bit combinations, $00,01,10$, and $11$, where the first and second bits respectively stand for $\text{U}_{1}$ and $\text{U}_{2}$ bits for a given bit duration.
Here, we assume that the total transmission power is fixed to $P$, that is, $E[(s_{BS}^{n})^{2}]=P$. Similar to the $\beta$ parameter of the uplink scheme, we define a new parameter, $\psi$, as the portion of $\text{U}_{1}$’s allocated power, that is, $m_{1,l}^{2}=m_{1,h}^{2}=\psi P$ (dc power of the transmitted signal). Accordingly, for $\text{U}_{2}$, we have $(\sigma_{2,l}^{2}+\sigma_{2,h}^{2})/2=(1-\psi)P$ (ac power of the transmitted signal). We again consider $m_{1,h}=-m_{1,l}$ and $\sigma_{2,h}^{2}=\alpha\sigma_{2,l}^{2}$.
@@
3.2 Downlink - User 1 Detection
The receiver architecture of the downlink scheme for $\text{U}_{1}$ is very similar to that of the uplink scheme, and the BEP can be evaluated by simple modifications. Specifically, $\text{U}_{1}$ calculates the sample mean of its received samples as $\bar{y}_{1}=\frac{1}{N}\sum\nolimits_{n=1}^{N}y^{n}_{1}$, which is also Gaussian distributed with $\mathrm{E}[\bar{y}_{1}]=h_{1}m_{1,i}$ and $\mathrm{VAR}[\bar{y}_{1}]=(|h_{1}|^{2}\sigma_{2,k}^{2}+\sigma_{w}^{2})/N$ for $i,k\in\{l,h\}$. Accordingly, the minimum distance detector is formulated as